# Quantum Learning and Estimation for Distribution Networks and Energy Communities Coordination

### v1

@@ Abstract
Price signals from distribution networks (DNs) guide energy communities (ECs) to adjust energy usage, enabling effective coordination for reliable power system operation. However, this coordination faces significant challenges due to the limited availability of information (i.e., only the aggregated energy usage of ECs is available to DNs), and the high computational burden of accounting for uncertainties and the associated risks through numerous scenarios. To address these challenges, we propose a quantum learning and estimation approach to enhance coordination between DNs and ECs. Specifically, leveraging advanced quantum properties such as quantum superposition and entanglement, we develop a hybrid quantum temporal convolutional network-long short-term memory (Q-TCN-LSTM) model to establish an end-to-end mapping between ECs’ responses and the price incentives from DNs. Moreover, we develop a quantum estimation method based on quantum amplitude estimation (QAE) and two phase-rotation circuits to significantly accelerate the optimization process under numerous uncertainty scenarios. Numerical experiments demonstrate that, compared to classical neural networks, the proposed Q-TCN-LSTM model improves the mapping accuracy by 69.2% while reducing the model size by 99.75% and the computation time by 93.9%. Compared to classical Monte Carlo simulation, QAE achieves comparable accuracy with a dramatic reduction in computational time (up to 99.99%) and requires significantly fewer computational resources.
Index Terms:
Quantum neural network, quantum amplitude estimation, energy communities, distribution networks, coordination
@@ I Introduction
With the rapid development of distributed energy resources (DERs), traditional centralized power systems are gradually evolving toward a more distributed architecture. Energy communities (ECs), which integrate various types of DERs, are playing a vital role in enhancing the local supply-demand balance and improving energy utilization efficiency [1]. In this context, achieving coordinated operation between distribution networks (DNs) and ECs is essential for enhancing system flexibility and reducing risks [2]. To this end, leveraging price incentives provided by DNs to generate grid-supportive responses of ECs is one of the effective approaches to facilitate such coordination.
@@
However, two critical challenges remain in achieving effective DNs-ECs coordination. The first challenge is that accurate estimation of ECs’ electricity consumption behavior remains difficult [3]. In some literature, an ideal assumption is made that the DNs have full access to ECs’ internal data (e.g., device parameters, operational states, and operational preferences). Thus, detailed operational models of ECs can be constructed based on physical modeling of devices, such as modeling electro-thermal coupling of air conditioners [4] and industrial production processes [5]. Moreover, to improve the overall computational tractability, simplified convex models were built and embedded into the centralized optimization model within the DNs [6]. However, in real-world situations, user electricity consumption behaviors usually exhibit strong nonlinearity, non-stationarity, and partial irrationality [7], making convex models inaccurate. Moreover, ECs are often unwilling or unable to share the device-level data due to privacy concerns, cybersecurity risks, or commercial interests [8], making the development of detailed user-side models impractical. Under such restrictions, DNs often can only access the aggregated energy usage of ECs, without insight into detailed device configurations or usage patterns [9]. To address this issue, distributed optimization methods (e.g., alternating direction method of multipliers [10]) have been widely adopted, where the centralized optimization problem is decomposed into several subproblems that are solved independently. While such methods offer good privacy protection, they often suffer from heavy communication burdens and slow convergence. Alternatively, some studies have proposed data-driven approaches to construct surrogate models that emulate user electricity consumption behaviors in response to external incentives [11]. However, such behavioral modeling is essentially a time-series analysis problem characterized by high dimensionality, temporal variability, and strong nonlinearity, leading to high learning complexity. In particular, for ECs with complex operational characteristics, achieving high behavioral modeling accuracy often requires a large number of model parameters [12], which further increases model complexity and reduces computational efficiency.
@@
The second challenge arises from the reliance on a large number of scenarios to characterize uncertainty-related risks [13] when precise probabilistic distributions are unavailable. For instance, conditional value-at-risk (CVaR)-based risk models require an accurate representation of the tail of distributions [14], which necessitates a large set of scenarios to capture the rare but high-impact events. Monte Carlo (MC) simulation is commonly employed to evaluate the expectation across a broad range of possible scenarios [15]. Nevertheless, MC methods often suffer from low computational efficiency and slow convergence. To reduce computational complexity, scenario reduction techniques have been explored to compress the original large-scale scenario set into a smaller yet representative subset [16]. However, for tail-sensitive risk measures such as CVaR, scenario reduction methods may not be able to fully preserve the integrity of the tail of distributions, leading to loss of valuable information and degraded risk awareness.
@@
Recent advances in quantum computing have opened up new opportunities for effective DNs-ECs coordination. Unlike classical computing, which processes information sequentially and requires large parameter sets to model complex systems, quantum computing leverages quantum superposition and entanglement properties inherent in quantum bits (qubits), allowing the representation and manipulation of an exponential number of states simultaneously with fewer parameters [17]. On the hardware front, companies such as IBM and Google have developed prototype systems with tens to hundreds of qubits, signaling the advent of the Noisy Intermediate-Scale Quantum (NISQ) era [18]. These developments make quantum computing particularly suitable for solving problems characterized by high dimensionality, complex data structures, and intensive computational demands with inherent parallelism.
@@
For the first challenge of estimating electricity consumption behavior of ECs, quantum neural networks (QNNs) offer a promising solution to establish the mapping between external incentives and ECs’ responses. Compared to classical neural networks, QNNs use qubits and quantum gates to perform computations in a superposed and entangled state, offering potential advantages in representing complex nonlinear relationships and accelerating the training process. Existing studies have applied QNNs to various power system-related tasks, such as wind speed forecasting [19], carbon price forecasting [20] and fault diagnosis [21]. However, the ECs incentive-response mapping problem incorporates ECs’ operational characteristics and constitutes a more complex temporal regression task. Simplified QNNs in current studies lack the capacity to sufficiently capture the intricate temporal and coupled dependencies inherent to ECs’ operation.
@@
For the second challenge of computational bottlenecks arising from large-scale scenario analysis, quantum estimation methods (e.g., quantum amplitude estimation, QAE) provide a promising solution. Compared to classical MC methods, QAE theoretically achieves a quadratic speedup [22, 23], demonstrating significant advantages in key subroutines such as expectation evaluation and CVaR computation [24]. Refs [25, 26] have applied QAE to power system reliability assessment, validating its potential in uncertainty-aware analysis. Most current research has focused on the theoretical feasibility of QAE algorithms, with limited discussions on the practical quantum circuit design. Therefore, a key challenge lies in designing task-specific quantum circuits that balance estimation accuracy and computational complexity to enable rapid and precise evaluation of uncertainty scenarios.
@@
    We develop a hybrid quantum temporal convolutional network–long short-term memory (Q-TCN-LSTM) model to estimate the energy consumption behavior of ECs in response to the price incentives from the DNs. To the best of our knowledge, the Q-TCN-LSTM model has not been explored in prior research. Q-TCN and Q-LSTM respectively extract short-term and long-term temporal dependencies, comprehensively modeling the coupled temporal dependencies of incentive-response characteristics of ECs. Compared to the classical TCN-LSTM, Q-TCN-LSTM improves accuracy by 69.2%, while reducing model size by 99.75% and computation time by 93.9%.
@@
    We develop a quantum amplitude estimation method to address the computational burden of estimating the gradient of CVaR under a large number of uncertainty scenarios, which is a critical component in DNs–ECs coordination. Two phase rotation circuits are introduced: one establishes analytical linear approximations between rotation angles and quantum states, achieving low circuit complexity but with potential approximation errors; the other computes exact rotation angles which are adaptively applied by introducing additional qubits, ensuring high accuracy but with increased complexity. Compared to the classical MC simulation, the proposed quantum estimation method achieves comparable accuracy with a dramatic reduction in computational time (up to 99.99%).
The rest of the paper is organized as follows. Section II introduces the coordination problem formulation between DNs and ECs. The methodology of the proposed quantum learning and estimation approach is presented in Section III. Numerical case studies are provided in Section IV. Finally, we summarize our conclusion in Section V.
@@ II Problem Formulation
II-A DNs-ECs Coordination Structure
The structure of DNs integrated with multi-type ECs and DERs (e.g., wind turbines (WT), photovoltaic (PV) systems, and new-type loads) is illustrated in Fig. 1. This paper focuses on three representative types of ECs: residential, commercial, and industrial.
Figure 1: Coordination structure between DNs and multi-type ECs.
In this paper, we focus on the day-ahead price-incentive demand response mechanism as a coordination strategy, wherein price signals guide ECs to adjust their operations accordingly. The coordination between DNs and various ECs is naturally a bilevel optimization problem, with DNs making decisions at the upper level and ECs responding at the lower level. At the upper level, DNs should accurately estimate the uncertainty associated with renewable energy sources (RESs), loads, and the trading power of ECs, and accordingly make risk-averse operational decisions, which is mainly considered as the price signals for ECs. At the lower level, each EC independently optimizes its operational strategy based on customized operational objectives in response to the price signals from the DNs, and submits the resulting power trading plans to the DNs.
@@ II-B Risk-Averse Operation Model of Distribution Networks
Considering the uncertainty in load and RESs, DNs need to make risk-averse decisions to reduce the operational risks including voltage violations and line overloads. In this paper, CVaR is adopted to quantify operational risk, where the uncertainties in load and RESs are modeled using a set of discrete scenarios. Accordingly, the overall operation model of the DNs is formulated as follows:
$\min$ $C=\sum_{s=1}^{S}\gamma_{s}f(z,\xi_{s})+\lambda(v_{\alpha}+\frac{1}{1-\alpha}\sum_{s=1}^{S}\gamma_{s}[f(z,\xi_{s})-v_{\alpha}]_{+}),$ (1a)
s.t. $f(z,\xi_{s})=\sum_{t=1}^{T}\Delta t[-\sum_{i\in\Omega^{\mathrm{B}}}\pi^{\mathrm{EC}}_{t}P^{\mathrm{EX,EC}}_{i,t}+\pi^{\mathrm{DN}}_{t}P^{\mathrm{EX,DN}}_{s,t}$
$\qquad\qquad+\beta_{1}\sum_{i\in\Omega^{\mathrm{B}}}([V_{i,s}-\overline{V}]_{+}+[\underline{V}-V_{i,s}]_{+})$
$\qquad\qquad+\beta_{2}\sum_{ij\in\Omega^{\mathrm{L}}}(P_{ij,s,t}-\overline{P}_{ij})],$ (1b)
@@
where (1c)-(1e) are the linearized distflow model. (1f)-(1g) are the power injection constraints. (1i)-(1j) are the price signal constraints. $\xi_{s}$ is the $s$-th scenario ($S$ in total) with probability $\gamma_{s}$. $f(z,\xi_{s})$ is the cost under the $s$-th scenario, where $z$ denotes the decision variables of DNs. $P^{\mathrm{EX,EC}}_{i,t}$ is the active trading power between the EC at node $i$ and DNs, which is the EC’s response to the price signal $\pi^{\mathrm{EC}}_{t}$. Note that the DNs do not know the exact operation patterns and device parameters of ECs, we adopt a unknown function $\phi_{i}(\cdot)$ to represent the corresponding response characteristics, as in (1h). The active and reactive trading power of ECs are assumed to be proportional, with the fixed ratio given by $P^{\mathrm{EX,EC}}_{i,t}/Q^{\mathrm{EX,EC}}_{i,t}=\kappa_{i}$. $\lambda$ is the risk aversion coefficient which controls the risk acceptance level. $v_{\alpha}$ is the VaR at the $\alpha$-th percentile. $[z]_{+}=\max(0,z)$. $\pi^{\mathrm{DN}}_{t}$ is the trading price published by the transmission network to DNs. $P^{\mathrm{EX,DN}}_{s,t}$ is the trading power of the DNs with the transmission network under scenario $s$ at time $t$. $\pi^{\mathrm{EC}}_{t}$ is limited by $\overline{\pi}^{\mathrm{EC}}_{t}/\underline{\pi}^{\mathrm{EC}}_{t}$, and averaged by $\pi^{\mathrm{EC,av}}_{t}$. $\beta_{1}/\beta_{2}$ are the penalty coefficients of the voltage and line power flow violations. $V_{i,s}$ is the voltage magnitude limited by $\overline{V}/\underline{V}$. $P_{ij,s,t}/Q_{ij,s,t}$ are the active/reactive power flow of line $ij$. $r_{ij}/x_{ij}$ are the resistance/reactance. $p_{j,s,t}/q_{j,s,t}$ are the active/reactive power injection. $P^{\mathrm{R}}_{j,s,t}$ is the active power output of RESs at bus $j$ under scenario $s$ at time $t$. $P^{\mathrm{L}}_{j,s,t}/Q^{\mathrm{L}}_{j,s,t}$ are the active/reactive power loads. In (1f), $P^{\mathrm{R}}_{j,s,t}=0$ and $P^{\mathrm{EX,EC}}_{j,s,t}=0$ if bus $j$ is not connected to any RES or EC, respectively.
II-C Incentive-Response Mapping Integrated Gradient Descent for DNs-ECs Coordination

### v2

# arXiv:2506.11730v2[math.OC] 27 Feb 2026

@@ Abstract
Price signals from distribution networks (DNs) guide energy communities (ECs) in adjusting their energy usage, enabling effective coordination for reliable power system operation. However, this coordinated operation faces significant challenges due to the limited availability of ECs’ internal information (i.e., only the aggregated energy usage of ECs is available to DNs), and the high computational burden of accounting for uncertainties and the associated risks through numerous scenarios. To address these challenges, we propose a quantum learning and estimation approach to enhance coordinated operation between DNs and ECs. Specifically, by leveraging advanced quantum properties such as quantum superposition and entanglement, we develop a hybrid quantum temporal convolutional network-long short-term memory (Q-TCN-LSTM) model to establish an end-to-end mapping between ECs’ responses and the price incentives from DNs. Moreover, we develop a quantum estimation method based on quantum amplitude estimation (QAE) and two phase-rotation circuits to significantly accelerate the optimization process under numerous uncertainty scenarios. Numerical experiments demonstrate that, compared to classical neural networks, the proposed Q-TCN-LSTM model improves the mapping accuracy by 69.2% while reducing the model size by 99.75%. Compared to classical Monte Carlo simulation, QAE achieves comparable accuracy with a substantial reduction in computational resources. In addition, the estimated computation time for quantum learning and estimation on ideal quantum devices is over 90% shorter than that of traditional methods.
@@ I Introduction
With the rapid development of distributed energy resources (DERs), traditional centralized power systems are gradually evolving toward a more decentralized architecture [1, 2]. Energy communities (ECs), which integrate various types of DERs, facilitate local consumption of renewable energy and enhance the local energy balance [3]. However, from the standpoint of upper-level distribution networks (DNs), the operational behaviors of ECs inherently exhibit uncertainties. In this context, achieving coordinated operation between DNs and ECs is essential for enhancing system flexibility and reducing potential risks (e.g., voltage issues, congestion) [4]. Among existing coordination methods, one effective and commonly used approach is to leverage price signals from DNs to guide grid-supportive responses from ECs [5]. However, two critical challenges remain in achieving effective DNs-ECs coordinated operation.
@@
The first challenge lies in accurately estimating ECs’ electricity consumption behavior [6, 7]. In the existing literature, an ideal assumption is often adopted that DNs possess full access to the internal operational data of ECs, including detailed device parameters and operational preferences. Consequently, precise operational models for ECs can be developed based on the physical characteristics of individual devices, such as the electro-thermal coupling dynamics of air conditioners [8] and specific industrial production processes [9]. Moreover, to improve the overall computational tractability, simplified convex EC operation models are built and embedded into the DNs operation model to form a centralized coordinated operation model [10]. However, this ideal assumption rarely holds true in practice for two primary reasons. First, user electricity consumption behaviors usually exhibit strong nonlinearity, non-stationarity, and partial irrationality, which cannot be accurately captured by simplified convex operational models. Second, ECs are unwilling or unable to share detailed, device-level data due to concerns about user privacy, cybersecurity risks, or commercial confidentiality [11], making it impractical to develop precise operational models based on internal data. Consequently, in practical applications, distribution networks usually have access only to aggregated energy consumption data from ECs without detailed insights into device-level configurations or usage patterns [12]. To address this issue, distributed optimization methods (e.g., alternating direction method of multipliers [13]) have been widely adopted, where the centralized optimization problem is decomposed into several subproblems that are solved independently. While such methods offer good privacy protection, they often suffer from heavy communication burdens and slow convergence. Alternatively, some studies have proposed data-driven approaches to construct surrogate models that emulate user electricity consumption behaviors in response to external incentives [14]. However, such behavioral modeling [15] is essentially a time-series analysis problem characterized by high dimensionality, temporal coupling, and strong nonlinearity, leading to high learning complexity. In particular, for ECs with complex operational characteristics, achieving high behavioral modeling accuracy often requires a large number of model parameters [16], which further increases model complexity and reduces computational efficiency.
@@
The second challenge arises from the substantial computational burden imposed by risk-aware models that rely heavily on a large number of discrete scenarios to accurately characterize uncertainties and potential operational risks, especially when precise probability distributions are unavailable [17]. This issue becomes especially pronounced when addressing extreme loss events using risk measures such as conditional value-at-risk (CVaR), which require accurate representations of distribution tails, thus demanding extensive scenario sampling to capture rare but high-impact events [18]. Typically, Monte Carlo (MC) simulation is employed to evaluate the risk expectation across a broad range of possible scenarios [19]. Nevertheless, MC’s estimation accuracy is correlated with the number of sample scenarios and often suffers from low computational efficiency and slow convergence. Scenario reduction techniques have been explored to compress the original large-scale scenario set into a smaller yet representative subset [20] to reduce computational complexity. However, for tail-sensitive risk measures such as CVaR, scenario reduction may not be able to fully preserve the integrity of distribution tails, leading to a loss of critical information and degraded risk-awareness.
@@
Recent advances in quantum computing have opened up new opportunities for effective DNs-ECs coordinated operation. Unlike classical computing, which processes information sequentially and requires large parameter sets to model complex systems, quantum computing leverages quantum superposition and entanglement properties inherent in quantum bits (qubits), allowing the representation and manipulation of an exponential number of states simultaneously with fewer parameters. On the hardware front, companies such as IBM and Google have developed prototype systems with tens to hundreds of qubits, signaling the advent of the Noisy Intermediate-Scale Quantum (NISQ) era [21]. These developments make quantum computing particularly suitable for solving problems characterized by high dimensionality, complex data structures, and intensive computational demands with inherent parallelism.
@@
To address the first challenge regarding the accurate estimation of ECs’ electricity consumption behavior, quantum neural networks (QNNs) present a promising approach for modeling the mapping between external incentives and EC responses. Compared to classical neural networks [22], QNNs leverage qubits and quantum gates to perform computations within superposed and entangled quantum states, potentially enhancing their capability to represent complex nonlinear relationships and accelerating the training process. Existing studies have demonstrated the effectiveness of QNNs in various power system applications, including carbon price forecasting [23] and fault diagnosis [24]. However, the incentive-response modeling of ECs involves heterogeneous operational characteristics, forming a significantly more intricate temporal regression task. The QNN architectures employed in current research remain relatively simplified, limiting their ability to fully capture the complex temporal dynamics and coupled operational dependencies inherent in EC behaviors.
@@
For the second challenge of computational bottlenecks arising from large-scale scenario evaluation, quantum estimation methods (e.g., quantum amplitude estimation, QAE) provide a promising solution. Compared to classical MC methods, QAE theoretically achieves a quadratic speedup [25] under specific conditions (e.g., efficient quantum state preparation and negligible quantum noise), demonstrating significant advantages in key subroutines such as expectation evaluation and CVaR computation [26]. Refs [27, 28] have applied QAE to power system reliability assessment, validating its potential in uncertainty-aware analysis. However, existing research predominantly addresses theoretical aspects of QAE algorithms, with limited exploration into practical quantum circuit implementations. Consequently, a significant challenge is designing quantum circuits capable of effectively balancing estimation accuracy and computational complexity, thus enabling rapid and precise uncertainty scenario evaluation in practical applications.
@@
    We develop a quantum temporal convolutional network–long short-term memory (Q-TCN-LSTM) model to estimate the energy consumption behavior of ECs in response to the price incentives from the DNs. To the best of our knowledge, the Q-TCN-LSTM model has not been explored in prior research. Q-TCN and Q-LSTM respectively extract short-term and long-term temporal dependencies, comprehensively modeling the coupled temporal dependencies of incentive-response characteristics of ECs. Compared to the classical TCN-LSTM, Q-TCN-LSTM improves accuracy by 69.2%, while reducing the model size by 99.75%. Moreover, on ideal quantum devices, the theoretically estimated learning time for one forward pass is over 90% shorter than that required by traditional learning methods.
@@
    We develop a quantum estimation method to address the computational burden of estimating the gradient of CVaR-based optimization under a large number of uncertainty scenarios, which is a critical component in DNs–ECs coordinated operation. Two phase rotation circuits are introduced: one establishes analytical linear approximations between rotation angles and quantum states, achieving low circuit complexity but with potential approximation errors; the other computes exact rotation angles which are adaptively applied by introducing additional qubits, ensuring high accuracy but with increased complexity. For the proposed quantum estimation method, the theoretically estimated computation time is reduced by up to 99.99% compared with classical Monte Carlo simulation.
The rest of the paper is organized as follows. Section II introduces the coordinated operation problem formulation between DNs and ECs. The methodology of the proposed quantum learning and estimation approach is presented in Section III. Numerical case studies are provided in Section IV. Finally, we summarize our conclusions in Section V.
@@ II Problem Formulation
In this section, we first present the DNs-ECs coordinated operation structure, followed by the detailed risk-averse operation model of DNs, and the coordination mechanism between DNs-ECs coordinated operation.
II-A DNs-ECs Coordinated Operation Structure
The structure of DNs integrated with multiple ECs and DERs (e.g., wind turbines (WTs), photovoltaic (PV) systems, and new-type loads) is illustrated in Fig. 1. This paper focuses on three representative types of ECs: residential (EC-R), commercial (EC-C), and industrial (EC-I). These three types not only cover the main categories of ECs in real-world situations thus ensuring the generalizability, but also exhibit significant heterogeneity across multiple dimensions (e.g., device composition, operational constraints and objectives, risk preferences), leading to distinct operational behaviors and introducing non-trivial complexity into DNs-ECs coordinated operation.
Figure 1: Coordinated operation structure between DNs and multiple ECs.
In this paper, we focus on the day-ahead price-incentive demand response mechanism as the coordinated operation strategy, where price signals guide ECs to adjust their operations accordingly. The coordinated operation between DNs and ECs is naturally a bilevel optimization problem, with DNs making decisions at the upper level and ECs responding at the lower level. At the upper level, DNs should accurately estimate the uncertainty associated with renewable energy sources (RESs), loads, and the aggregated energy consumption behaviors (i.e., trading power) of ECs, and accordingly determine risk-averse operational decisions, which is mainly considered as the price signals for ECs. At the lower level, each EC independently optimizes its operational strategy based on customized operational objectives in response to the price signals from the DNs, and submits the resulting power trading plans to the DNs.
@@ II-B Risk-Averse Operation Model of Distribution Networks
Considering the uncertainties in load and RESs, DNs need to make risk-averse decisions to reduce the operational risks including voltage violations and line overloads. In this paper, CVaR is adopted to quantify operational risks, where the uncertainties in load and RESs are modeled using a set of discrete scenarios. Accordingly, the overall operational model of the DNs is formulated as follows:
$\min_{\pi^{\mathrm{EC}},v_{\alpha}}\$$C=\sum_{s=1}^{S}\gamma_{s}C_{s}+\lambda(v_{\alpha}+\frac{1}{1-\alpha}\sum_{s=1}^{S}\gamma_{s}[C_{s}-v_{\alpha}]_{+}),$ (1a)
s.t. $C_{s}=\sum_{t=1}^{T}\Delta t[-\sum_{i\in\Omega^{\mathrm{EC}}}\pi^{\mathrm{EC}}_{t}P^{\mathrm{EX,EC}}_{i,t}+\pi^{\mathrm{DN}}_{t}P^{\mathrm{EX,DN}}_{s,t}$
$\quad+\beta_{1}\sum_{i\in\Omega^{\mathrm{B}}}([V_{i,s,t}-\overline{V}_{i}]_{+}+[\underline{V}_{i}-V_{i,s,t}]_{+})$ (1b)
$\quad+\beta_{2}\sum_{ij\in\Omega^{\mathrm{L}}}([P_{ij,s,t}-\overline{P}_{ij}])_{+}+[\underline{P}_{ij}-P_{ij,s,t}]_{+})],$
@@
where the released incentive price $\pi^{\mathrm{EC}}$ and the VaR $v_{\alpha}$ are the decision variables of DNs. (1c)-(1e) are the linearized distflow model. (1f)-(1g) are the power injection constraints. (1i)-(1j) are the price signal constraints. $C_{s}$ is the cost under the $s$-th scenario ($S$ in total) of uncertain DERs integrated to the DNs with probability $\gamma_{s}$. $P^{\mathrm{EX,EC}}_{i}$ is the active trading power between the EC at node $i$ and DNs, which is the EC’s response to the price signal $\pi^{\mathrm{EC}}$. Note that the DNs do not know the exact operation patterns and device parameters of ECs, we adopt an unknown function $\phi_{i}(\cdot)$ to represent the corresponding response characteristics, as in (1h). The active and reactive trading power of ECs are assumed to be proportional. $\lambda$ is the risk aversion coefficient which controls the risk acceptance level. $z]_{+}=\max(0,z)$. $\pi^{\mathrm{DN}}_{t}$ is the trading price published by the transmission network to DNs. $P^{\mathrm{EX,DN}}_{s,t}$ is the trading power of the DNs with the transmission network under scenario $s$ at time $t$. $\pi^{\mathrm{EC}}_{t}$ is bounded by $\overline{\pi}^{\mathrm{EC}}_{t}/\underline{\pi}^{\mathrm{EC}}_{t}$, and averaged by $\pi^{\mathrm{EC,av}}_{t}$. $\beta_{1}/\beta_{2}$ are the penalty coefficients of the voltage and line power flow violations. $V_{i,s}$ is the voltage magnitude bounded by $\overline{V}_{i}/\underline{V}_{i}$. $P_{ij,s,t}/Q_{ij,s,t}$ are the active/reactive power flow of line $ij$. $r_{ij}/x_{ij}$ are the resistance/reactance. $p_{j,s,t}/q_{j,s,t}$ are the active/reactive power injection. $P^{\mathrm{R}}_{j,s,t}$ is the active power output of RESs at bus $j$ under scenario $s$ at time $t$. $P^{\mathrm{L}}_{j,s,t}/Q^{\mathrm{L}}_{j,s,t}$ are the active/reactive power loads. $\Omega^{\mathrm{EC}}/\Omega^{\mathrm{B}}/\Omega^{\mathrm{L}}$ denote the sets of nodes with EC, nodes, and lines, respectively. In (1f), $P^{\mathrm{R}}_{j,s,t}=0$ and $P^{\mathrm{EX,EC}}_{j,s,t}=0$ if bus $j$ is not connected to any RES or EC, respectively.
II-C DNs-ECs Coordinated Operation